In a race , Mary is 50 ft. in front of Susie after 10 seconds. How fast can Mary run if Susie runs 20 ft per second?
Answer by bucky(2097)

The distance formula tells that Distance = Rate * Time.
.
If Susie runs at 20 ft per second, then in 10 seconds she has run 200 ft. You can get this
from the Distance formula:
.
Distance = 20 * 10 = 200
.
But in that same 10 seconds Mary has run 50 feet more than Susie. So in 10 seconds Mary has run
250 feet.
.
Now you can calculate Mary's Rate (which is her speed). You know the distance she ran was
250 feet and the time that she ran it in was 10 seconds. Substituting those values into the
distance formula gives you:
.
250 = R * 10
.
Solve for R by dividing both sides of this equation by 10 to get:
.
R = 250/10 = 25 feet per second.
.
So while Susie runs at a rate of 20 ft per second, Mary (who is faster) runs at a rate of
25 feet per second.
.
Hope this helps you to understand the problem.
.